The mean and standard deviation of a random sample of n measurements are equal to 34.1 and 3.5​, respectively. a. Find a 90​% confidence interval for mu if nequals64. b. Find a 90​% confidence interval for mu if nequals256. c. Find the widths of the confidence intervals found in parts a and b. What is the effect on the width of a confidence interval of quadrupling the sample size while holding the confidence coefficient​ fixed?
Answer #1

a)

sample mean sample size population std deviation 34.10 3.500 邨: standard errror of mean= 0%-o/vn- 0.4375 for 90 % CI value of

b)

standard errror of mean = σx=σ/√n= 0.2188
for 90 % CI value of z= 1.645
margin of error E=z*std error                            = 0.360
lower confidence bound=sample mean-margin of error= 33.74
Upper confidence bound=sample mean +margin of error= 34.46

c)

width got halved quadrupling the sample size